Find the amount to be paid at the end of 3 years in each case:  Principal = ₹ 1,200 at 12% p.a. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given: - Principal (P)= ₹1200, Rate (R) = 12% p.a. and Time (T) = 3years.
If interest is calculated uniformly on the original principal throughout the loan period, it is called Simple interest (SI).
SI = (P × R × T)/100
= (1200 × 12 × 3)/ 100
= (12 × 12 × 3)/ 1
= ₹432
Amount = (principal + SI)
= (1200 + 432)
= ₹1632

Thanks